Weekly Report 6-20-2012

Categories: Everything else Circ

User Services Weekly Report
Date: 6-20-2012
Name: Dan Frostman
Circulation Services – Circ, Reserves, and Stacks

Statistics:
• Number of Google searches so far today: 1,425,758,645 (according to http://www.worldometers.info/)
Highlights: (Progress towards Goals, accomplishments)
Armstrong:
• Very quiet
• Staff on vacation
• Planning to get new plotter installed this week
• Working on Biological abstracts
• Pulling weeding lists
Davis Family:
• Trying to keep the students busy
• Work on the self-check machine (still dead)
• A group on campus is investigating the possibility of circulating bike lock keys at circ
• Goals setting
• Gave lesson to prospective oboe student
• HEAT replacement narrowed to two vendors – will set up demos
• Set up bi-weekly meetings for all circ staff
Stacks:
• Cleaning
• Shifting
Reserves:
• Packing up BL reserves
• LS reserves

Security team: encrypted email demos
Space Team: n/a
CT Team: Segue migration workshops prep

Challenges:
• Still waiting on BL reserve recalls

Leave a Reply